A 100 KG WOODEN CRATE IS AT REST ON A LEVEL STONE FLOOR WITH COEFFICIENT OF FRICTION OF 0.5 .

A. WHAT IS THE MINIMUM HORIZONTAL FORCE NEEDED TO START THE CRATE MOVING ?

B. WHAT IS THE MINIMUM HORIZONTAL FORCE NEEDED TO KEEP THE CRATE IN MOTION AT CONSTANT SPEED WHERE THE COEFFICIENT OF KINETIC FRICTION IS 0.4 ?


1
Expert's answer
2022-02-20T15:43:23-0500

Given:

"m=100\:\rm kg"

"\mu_s=0.5,\quad \mu_k=0.4"


A. WHAT IS THE MINIMUM HORIZONTAL FORCE NEEDED TO START THE CRATE MOVING ?

"F=\mu_sN=\mu_smg=0.5*9.8*100=490\:\rm N"

B. WHAT IS THE MINIMUM HORIZONTAL FORCE NEEDED TO KEEP THE CRATE IN MOTION AT CONSTANT SPEED WHERE THE COEFFICIENT OF KINETIC FRICTION IS 0.4 ?

"F=\mu_kN=\mu_kmg=0.4*9.8*100=392\:\rm N"